# TrojanForge: Adversarial Hardware Trojan Examples with Reinforcement Learning

### v1

@@ 4.1 JSI and Trigger Compatibility
Table 2 shows Q-coverage (explained in Section 3.2) for the combinational ISCAS-85 benchmark circuits (listed in each row) when using random test vectors, $D1$, $D2$, $D3$ from Sarihi et al. (2023), and DETERRENT (labeled in each column, respectively).
@@
DETERRENT only includes test vectors for four circuits ($c2670,c5315,c6288,c7552$), and it provides higher Q-coverage than other detectors except for the case of $c2670$. While Gohil et al. Gohil et al. (2022b) shows 100% four-input coverage with 8 test vectors for $c2670$, this is not the case with our candidate selection strategy. Such interpretations could lead to false security interpretations and vulnerability exploitation by attackers. By selecting 20 candidate trigger nets ($\lvert M\lvert=20$), up to 15,504 checks are needed for each insertion by the detector; however, circuit structures and detectors’ imperfections lead to significantly fewer checks.

### v3

@@ 3.1 JSI and Trigger Compatibility
Table 2 shows Q-coverage (explained in Section 2.2) for the combinational ISCAS-85 benchmark circuits (listed in each row) when using random test vectors, $D1$, $D2$, $D3$ from [18], and DETERRENT (labeled in each column, respectively).
@@
DETERRENT only includes test vectors for four circuits ($c2670,c5315,c6288,c7552$), and it provides higher Q-coverage than other detectors except for the case of $c2670$. While Gohil et al. [11] shows 100% four-input coverage with 8 test vectors for $c2670$, this is not the case with our candidate selection strategy. Such interpretations could lead to false security interpretations and vulnerability exploitation by attackers. By selecting 20 candidate trigger nets ($\lvert M\lvert=20$), up to 15,504 checks are needed for each insertion by the detector; however, circuit structures and detectors’ imperfections lead to significantly fewer checks.
@@ 4 Conclusion
Acknowledgments
This work has been partially funded by NSF grants 2219680 and 2219679.